solve for the x given BD= 3x +3 and AE = 4x+7. Assume B is the midpoint of AC and D is the midpoint of CE
Answer by midgetwill(2)

If you were to draw this out, AE would be twice the length of BD, so you would start with:
2(3x+3) = 4x+7 - Which simplifies to:
6x+6 = 4x+7 - Subtract 4x and 6 from both sides:
2x = 1 - Divide 1 by 2:
x= 1/2